OpenStudy (highschoolmom2010):

A scientist discovered that a certain element was present in 35% of the samples she studied. In 15% of all samples, the element was found in a special compound. What is the probability that the compound is in a sample that contains the element?

OpenStudy (highschoolmom2010):

certain element was present in 35% of the samples =.35 15% of all samples, the element was found in a special compound=.15

ganeshie8 (ganeshie8):

we need to find P(compound | element)

ganeshie8 (ganeshie8):

P(A | B) = P(A and B)/P(B)

ganeshie8 (ganeshie8):

P(compound | element) = P(compound and element) / P(element) = 0.15/0.35 = 42.86 %

ganeshie8 (ganeshie8):

check if thats the correct answer .-.

OpenStudy (highschoolmom2010):

my book has the answer listed as 3/7

OpenStudy (akashdeepdeb):

That is 42.86%!

OpenStudy (akashdeepdeb):

3/7 = 42.86% [When it is represented as percentages. :) ]

OpenStudy (highschoolmom2010):

oh hehe i knew that